Business and finance. There are 850 Douglas fir and ponderosa pine trees in a
section of forest bought by Sawz Logging Co. The company paid an average of
$300 for each Douglas fir and $225 for each ponderosa pine. If the company paid
$217,500 for the trees, how many of each kind did the company buy?

Answer by ankor@dixie-net.com(22740)

There are 850 Douglas fir and ponderosa pine trees in a
section of forest bought by Sawz Logging Co. The company paid an average of
$300 for each Douglas fir and $225 for each ponderosa pine. If the company paid
$217,500 for the trees, how many of each kind did the company buy?
:
Let D = number of Douglas firs; P = number of Ponderosas
:
Total trees:
D + P = 850
or
D = 850 - P
:
Cost of all the trees:
300D + 225P = 217500
:
Substitute (850-P) for D, solve for P:
300(850-P) + 225P = 217500
:
255000 - 300P + 225P = 217500
:
-300P + 225P = 217500 - 255000
:
-75P = -37500
:
P = -37500/-75
:
P = 500 Ponderosa trees
:
850 - 500 = 350 Douglas firs.
:
:
Check:
225(500) + 300(350) =
112500 + 105000 = 217500